From the top of tower 100m high, a ball is dropped with an angle of 45^°. If it strikes 100m away from the tower then what will be displacement of the ball?

Grade:12th pass

2 Answers

Kunal Kumar
5 years ago
L
Rajdeep
5 years ago
HELLO THERE!
 
Consider the total thing that happened to be a right angled triangle. 
 
Let the tower be the height of the triangle, 
the displacement of the ball is the hypotenuse of the triangle (say, x)
the distance from the tower where the ball falls is the base of the triangle.
 
Given: Base and height = 100m
 
Now the angle which the base (distance of ball from tower, makes with the tower is 90o). Since angle between the hypotenuse (path of the ball) and height (height of tower) is 45o (given), the other angle between the hypotenuse and the base is also 45o.
 
So, we just have to find out the hypotenuse of the triangle to find out the displacement of the ball from the tower.
 
If the hypotenuse is x,
 
x = \sqrt{b^{2} + h^{2}} \\\\=\sqrt{100^{2} + 100^{2}} =100\sqrt{2}m
 
THANKS!